An alphabetical list of the adjusted trial balance accounts for North Country Rentals after its first year of operations ending March 31, 2020, is shown below:

Account Adjusted
Account Balance*
Accounts payable $ 9,600
Accumulated depreciation, building 25,500
Accumulated depreciation, furniture 4,000
Advertising expense 16,700
Building 596,000
Cash 17,500
Depreciation expense, building 25,500
Depreciation expense, furniture 4,000
Furniture 43,800
Interest expense 10,310
Interest payable 850
Janitorial expense 41,500
Land 115,000
Long-term notes payable 367,000
Notes receivable, due 2023 148,000
Office salaries expense 129,125
Office supplies 750
Office supplies expense 6,600
Brand name 3,500
Prepaid advertising 450
Rent revenue 403,400
Rent receivable 16,500
Salaries payable 2,875
Utilities expense 37,220
Wyett North, capital 427,730
Wyett North, withdrawals 28,500


*Assume all accounts have a normal balance.
Required:
1.
Calculate the capital balance as it would appear on the March 31, 2020, balance sheet.
2. Prepare a classified balance sheet. Assume that $217,500 of the Long-Term Notes Payable will be paid during the year ended March 31, 2021. Also, $57,500 of the notes receivable will be collected by March 31, 2021.
3. Calculate the current ratio and the debt to equity ratio. (Round the final answers to 2 decimal places.)
Answer #1
Part-1 : Computation of Capital Balance
Beg Balance $427,730.00
Net Income $132,445.00
Drawings $28,500.00
Ending Capital BAlance $531,675.00
Part-2 Balance Sheet - North Country Rentals
Amount Amount
Asset
Current Assets
Cash $17,500.00
Rent Receivable $16,500.00
Prepaid Advertising $450.00
Office Supplies $750.00
Total Current Assets $35,200.00
Investments
Notes Receivable, Due 2023 $148,000.00 $148,000.00
Property, Plant & Equipment
Land $115,000.00
Building $596,000.00
Accumulated Depreciation, Building -$25,500.00
Furniture $43,800.00
Accumulated Depreciation, Furniture -$4,000.00
Total Property, Plant & Equipment $725,300.00
Intangible Assets
Brand Name $3,500.00 $3,500.00
Total Assets $912,000.00
Liabilities
Current Liabilities
Accounts Payable $9,600.00
Interest Payable $850.00
Salaries Payable $2,875.00
Total Current Liabilities $13,325.00
Long Term Liabilities
Long Term Notes Payable $367,000.00 $367,000.00
Total Liabilities $380,325.00
Equity
Wyett Noth' Capital $531,675.00
Total Liabilities & Equity $912,000.00
Part-3
Current Ratio= Current Asset/ Current Liabilities
=35200/13325=2.64
Debt to Equity Ratio= Debt/ Equity
=380325/531675=0.72